Emily has factored a particular polynomial as (a-b)(x-y). George factors the same polynomial and gets (b-a)(y-x). Who is correct and why?

Answer by narayaba(40)

Both of them are correct
(a-b)(x-y) can be rewritten as follows
(a-b) = - (b-a) similarly (x-y)= -(y-x)
(a-b)(x-y) = -(b-a)*-(y-x) = -1*-1 (b-a)(y-x) = (b-a)(y-x)